# Lumos-1: On Autoregressive Video Generation from a Unified Model Perspective

### v1

@@ Abstract
Autoregressive large language models (LLMs) have unified a vast range of language tasks, inspiring preliminary efforts in autoregressive video generation. Existing autoregressive video generators either diverge from standard LLM architectures, depend on bulky external text encoders, or incur prohibitive latency due to next-token decoding. In this paper, we introduce Lumos-1, an autoregressive video generator that retains the LLM architecture with minimal architectural modifications. To inject spatiotemporal correlations in LLMs, we identify the efficacy of incorporating 3D RoPE and diagnose its imbalanced frequency spectrum ranges. Therefore, we propose MM-RoPE, a RoPE scheme that preserves the original textual RoPE while providing comprehensive frequency spectra and scaled 3D positions for modeling multimodal spatiotemporal data. Moreover, Lumos-1 resorts to a token dependency strategy that obeys intra-frame bidirectionality and inter-frame temporal causality. Based on this dependency strategy, we identify the issue of frame-wise loss imbalance caused by spatial information redundancy and solve it by proposing Autoregressive Discrete Diffusion Forcing (AR-DF). AR-DF introduces temporal tube masking during training with a compatible inference-time masking policy to avoid quality degradation. By using memory-efficient training techniques, we pre-train Lumos-1 on only 48 GPUs, achieving performance comparable to EMU3 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V. Code and models are available at https://github.com/alibaba-damo-academy/Lumos.
@@ 1 Introduction
Autoregressive models have demonstrated significant advancement in the field of language processing [6, 44, 70, 45, 3, 37] by unifying diverse language tasks into one single framework and expanding the scale of large language models to unprecedented sizes. Following this trail of research, significant efforts have emerged in autoregressive visual generation [33, 7, 8, 42]. By utilizing similar architectural and generation designs as those used in LLMs, there is substantial potential for advancing LLMs towards a unified model capable of both visual generation and understanding [77, 81, 95, 68].
@@
Pioneering research efforts have been dedicated to instantiating autoregressive visual generation, with a focus on key aspects, spanning improving autoregressive paradigm [46, 67, 8, 42], improving tokenizer format [33] (e.g., from discrete to continuous), and enhancing tokenizer capabilities [64, 1, 74, 58]. However, an autoregressive video generation paradigm that is fully compatible with unified models remains underexplored. Preliminary attempts either exhibit architectural distinctions with LLM architectures [14, 72], rely on external text encoders [14], or trail in generation efficiency [79], leaving much room for advancement.
In this work, we propose Lumos-1, a model that leverages the exact architecture of LLM with minimal architectural modifications to achieve autoregressive video generation, in such a way that it enables video generation without a pre-trained external text encoder (figure 1 visualizes examples generated by Lumos-1). To ensure inference efficiency, we build Lumos-1 upon the discrete diffusion paradigm [81, 63], whose lower bound is the cross-entropy loss employed in MaskGIT [8].
@@
Distinct from texts, videos inherently exhibit strong spatiotemporal correlation. Although previous works have incorporated this prior into generative models by overturning the next-token prediction paradigm [67, 8], we dive into a basic technique that’s widely used in LLMs—Rotary Position Embeddings (RoPE), which is heavily underexplored in autoregressive video generation. Preliminary experiments conducted with Lumos-1 reveal that 1D RoPE is far from optimal, and the incorporation of 3D RoPE facilitates autoregressive generative learning. However, existing 3D RoPE suffers from imbalanced frequency spectrum ranges for temporal and spatial modeling. Building on this insight, we propose MM-RoPE, an improved version of RoPE that better accommodates the structure of video data. This enhancement involves more distributed frequency allocations for more comprehensive context modeling and the appropriate scaling of 3D position information for modality balancing. Though MM-RoPE modifies RoPE for visual tokens, the RoPE for text tokens remains consistent with the one from original LLMs, therefore preserving the language learning capabilities.
@@
Distinct from images, videos inherently exhibit 1) temporal causality between frames and 2) spatial information redundancy between frames. The temporal causality renders the paradigm of autoregressive image generation suboptimal for videos, which relies on global sequence bidirectional dependency [8] or raster-scan order dependency [62]. In response, Lumos-1 resorts to a token dependency strategy that emphasizes intra-frame bidirectional dependencies and inter-frame temporally causal dependencies, which resonates with the nature of video data. Meanwhile, spatial information redundancy degrades the efficacy of training with random mask prediction, as spatial information leakage allows the model to easily predict the masked token by attending to unmasked ones in previous frames. To alleviate this, Lumos-1 introduces Autoregressive Discrete Diffusion Forcing (AR-DF). The core of AR-DF training involves temporal tube masking [69], which repeats a frame-level random mask pattern across the temporal dimension. To some extent, such a design mitigates the issue of frame-wise imbalanced loss [79], thereby enabling more effective learning. The core of AR-DF inference includes an inference-time masking strategy that aligns with partial observation of history during training, enabling inference without degraded frame quality and motion.
@@
We leverage Llama [71] as our architecture and leverage a unified discrete codebook for language and visual data. Through a stage-wise training with GPU memory-friendly techniques, we pre-tarin Lumos-1 from scratch on 60 million images and 10 million videos using only 48 GPUs, achieving performance comparable to EMU3 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V.
@@ 2 Related Work
Autoregressive image generation. Spurred by the progress of LLMs, visual generation research is rapidly shifting from diffusion paradigms [55, 24, 93, 15] to an autoregressive one [62, 67, 86, 27, 52, 96, 90, 23, 43, 75, 53], due to its potential of being integrated into unified models [65, 97] with minimal modification. Preliminary research, such as Parti [87], DALL-E [51], MaskGiT [8] and LLamaGen [62], have demonstrated the efficacy of autoregressive generation using discrete tokenizers. MAR [33] incorporates a diffusion-style objective into autoregressive training, thereby accommodating continuous tokens. VAR [67] and FAR [86] reformulate the autoregressive target as next-scale and next-frequency prediction, respectively, improving generation fidelity. Unlike them, we dive into the design of the RoPE technique in visual generation.
Autoregressive video generation. Autoregressive video synthesis follows a trajectory similar to images, but the substantial computational overhead encouraged researchers to explore different granularities of autoregression. At the macro level, methods such as CausVid [84], Pyramidal Flow [29], and MAGI-1 [56] generate video clips by recursively invoking diffusion models. A second line of work adopts a hybrid AR-diffusion strategy that couples a pre-trained diffusion backbone with an external autoregressive planner, as exemplified by ARLON [34] and Mardini [40]. Our focus, however, is on micro AR approaches [21, 96, 88, 89, 72, 30, 77, 14, 79, 85], which treat the entire spatiotemporal token sequence as a single context and generate videos end-to-end with one autoregressive transformer (e.g., Phenaki [72], VideoPoet [30], Loong [79], EMU3 [77], NOVA [14]). Building on this paradigm, we introduce AR-DF, which is compatible with intra-frame bidirectionality and inter-frame temporal causality, enabling effective training and inference with LLM-style architectures.
RoPE for vision-language data. RoPE was first proven effective for LLMs [61, 70] and later adopted in DiTs [47] to inject spatial priors. However, its potential for autoregressive video generation remains heavily underexplored. Related works, including M-RoPE [76], RoPE-2D [2], TAD-RoPE [17] and VideoRoPE [80], incorporated structure priors (i.e., spatiotemporal information or its subsets) into RoPE for better vision-language understanding. VoPE [10] improves high-resolution image generation by continuous resolution learning. We propose MM-RoPE, a distributed and scaled RoPE technique that improves AR video synthesis while remaining plug-and-play for unified multimodal models.
@@ 3 Lumos-1
This section aims to introduce the design philosophy behind Lumos-1. In Sec. 3.1, we introduce MM-RoPE that enables better spatiotemporal awareness of an LLM for modeling visual data. In Sec. 3.2, we introduce AR-DF that enables effective training and inference. In Sec. 3.3, we introduce important techniques for implementing Lumos-1, including architecture, memory-friendly techniques, etc.
@@ 3.1 Spatiotemporal Correlation Injection via MM-RoPE
Preliminaries of 3D RoPE. One de facto design of contemporary LLM is the RoPE [61] technique, whose overall aim is to encode the absolute position with a rotation matrix while incorporating the explicit relative position dependency in the attention mechanism. This can be formulated as:
where $f_{q}(\bm{x}_{m},m)$ encodes the position $m$ for the embedding $\bm{x}_{m}$ to obtain the query feature ($f_{k}(\bm{x}_{n},n)$ is analogously defined); $g(\bm{x}_{m},\bm{x}_{n},m-n)$ is a function that defines the inner product between the query and key vectors that explicitly encodes the relative position $m-n$. The resultant form of function $f_{\{q,k\}}$ can be formulated as:
$f_{\{q,k\}}(\bm{x}_{m},m)=\bm{R}^{d}_{\Theta,m}\bm{W}_{q,k}\bm{x}_{m}$ (2)
$\begin{split}f_{q}(\bm{x}_{m},m)^{\rm T}f_{k}(\bm{x}_{n},n)&=\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}\bm{R}^{d}_{\Theta,\tau}\bm{W}_{k}\bm{x}_{n}\\ \tau&=n-m\end{split}$ (3)
$\begin{split}\bm{R}^{d}_{\Theta,\tau}&=\begin{bmatrix}R_{\theta_{1},\tau}&0&\cdots&0\\ 0&R_{\theta_{2},\tau}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\cdots&R_{\theta_{d/2},\tau}\\ \end{bmatrix}\\ R_{\theta,\tau}&=\begin{bmatrix}\cos\tau\theta&-\sin\tau\theta\\ \sin\tau\theta&\cos\tau\theta\end{bmatrix}\end{split}$ (4)
However, the application of the original RoPE to modeling visual data remains suboptimal considering the spatiotemporal correlation of visual tokens. One popular type of generative models, diffusion models [25, 59, 24], have improved upon this technique by proposing 3D RoPE that jointly injects spatiotemporal latent coordinates during attention calculation and demonstrates its effectiveness [26, 31, 66]. If we slightly abuse the annotation by denoting $\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}$ and $\bm{W}_{k}\bm{x}_{n}$ as $\bm{X}_{m}^{\rm T}$ and $\bm{X}_{n}$, we can formulate the attention calculation in equation 1 based on 3D RoPE as:
$\bm{X}_{m,t_{s}:t_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{t_{s}+1},\tau_{t}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{t_{e}},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,t_{s}:t_{e}}+$ (5)
$\bm{X}_{m,h_{s}:h_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{h_{s}+1},\tau_{h}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{h_{e}},\tau_{h}}\\ \end{bmatrix}\bm{X}_{n,h_{s}:h_{e}}+$
$\bm{X}_{m,w_{s}:w_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{w_{s}+1},\tau_{w}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{w_{e}},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,w_{s}:w_{e}}$
where $\{t_{s},t_{e}\}=\{0,\frac{2}{16}d\}$, $\{h_{s},h_{e}\}=\{\frac{2}{16}d,\frac{5}{16}d\}$ and $\{w_{s},w_{e}\}=\{\frac{5}{16}d,\frac{1}{2}d\}$ denote the start and end dimension index for encoding temporal, height and width relative position; $\bm{X}_{m,t_{s}:t_{e}}^{\rm T}$ denotes the submatrix extracted from $\bm{X}_{m}^{\rm T}$ using row indices $[2t_{s},2t_{e})$; other matrices are similarly defined.
@@
Initial exploration of incorporating 3D RoPE. Our preliminary exploration starts with the introduction of 3D RoPE to autoregressive video generation. We follow previous works [16, 33] to utilize the validation loss to observe the effectiveness due to its strong correlation with evaluation metrics. By default, we use the cross-entropy loss (C-Loss), following the standard LLM training objective. We compare the vanilla LLM RoPE with three schemes, as shown in figure 2: 1) Scheme 1, which allocates the first $1/2$ channels to encode the global position (i.e., the index in the global sequence) and the second $1/2$ channels to encode the temporal, height and width positions with a ratio of $2:3:3$. For text tokens, we only use the first half channels to encode the global positions to ensure the language modeling capability, while for visual tokens, we only use the second half to encode 3D positions; 2) Scheme 2, which extends upon scheme 1 by leveraging the first half channels of the visual tokens to encode the global positions. 3) M-RoPE [76], which uses all channels of the visual tokens to encode the 3D positions. From figure 3(a), we can observe that: 1) The incorporation of the spatiotemporal correlation in RoPE can significantly improve the performance of the model fitting data by comparing vanilla RoPE and Scheme 1. 2) The injection of the raster-scan order position information to visual tokens (i.e., the global position in Scheme 2 of figure 2(a)) can degrade the performance. 3) A comprehensive channel utilization (M-RoPE) is better than partial channel utilization (Scheme 1). Therefore, it is promising to inject such priors in this generative model.
@@
Peering into 3D RoPE and identifying its limitations. Although 3D RoPE proves effective in practice, its design remains suboptimal. In figure 3(b), we visualize how the frequencies are allocated to model temporal, height, and width dimensions. We observe that the temporal channels dominate large frequency spectrum, whereas the height and width channels are relegated to near-zero frequencies. For the sine function, the relative positions $\tau$ (when $\tau>=0$) should not exceed one period to avoid ambiguity since radians beyond $2\pi$ start repeating patterns in the function. Beyond this range, the model cannot distinguish fine-grained positional differences since rotary embeddings of this channel lose uniqueness. For the low-indexed channels, the embeddings can rotate significantly faster than high-indexed channels (as shown in figure 3(c)), leading to accelerated aliasing and loss of embedding uniqueness. For the high-indexed channels, the embeddings rotate so slowly that they lack sufficient resolution to model subtle local changes. Finally, while height and width are symmetrically important, they occupy disproportionately small and different segments of the overall frequency spectrum, reducing their capacity to capture spatial details effectively.
@@
MM-RoPE: a distributed and scaled 3D RoPE mechanism. To elegantly solve the aforementioned limitation, we propose MM-RoPE, a distributed 3D RoPE mechanism. Compared with M-RoPE [76] that is widely adopted in vision-language models, one core idea of MM-RoPE is to encode relative positions across comprehensive frequency spectra for all 3D information. As illustrated in figure 2(b), the RoPE for text tokens in MM-RoPE follows the standard design of an LLM, whereas the RoPE for visual tokens is comprised of multiple meta MM-RoPE components. Within each meta MM-RoPE, we keep the ratio of 3D information identical to 3D RoPE (i.e., $2:3:3$), while minimizing the overall dimensionality to maintain a more distributed design. Concretely, we first allocate channels for temporal information, then symmetrically interleave height and width channels to model spatial information. We can formulate the attention calculation of the first meta MM-RoPE as:
$\bm{X}_{m,0:2}^{\rm T}\begin{bmatrix}R_{\theta_{1},\tau_{t}}&0\\ 0&R_{\theta_{2},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,0:2}+$ (6)
$\bm{X}_{m,2:4}^{\rm T}\begin{bmatrix}R_{\theta_{3},\tau_{h}}&0\\ 0&R_{\theta_{4},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,2:4}+$
$\bm{X}_{m,4:6}^{\rm T}\begin{bmatrix}R_{\theta_{5},\tau_{h}}&0\\ 0&R_{\theta_{6},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,4:6}+$
$\bm{X}_{m,6:8}^{\rm T}\begin{bmatrix}R_{\theta_{7},\tau_{h}}&0\\ 0&R_{\theta_{8},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,6:8}$
@@
3:$\mathcal{C}\leftarrow G_{\phi}(\bm{X}_{p},AttnMask^{(p)})$triangleright Store cache for the prompt

### v2

# Lumos-1: On Autoregressive Video Generation with Discrete Diffusion from a Unified Model Perspective

@@ Abstract
Autoregressive large language models (LLMs) have unified a vast range of language tasks, inspiring preliminary efforts in autoregressive (AR) video generation. Existing AR video generators either diverge from standard LLM architectures, depend on bulky external text encoders, or incur prohibitive latency due to next-token decoding. In this paper, we introduce Lumos-1, an LLM-based unified model for AR video generation with efficient discrete diffusion. Firstly, to fit videos with LLMs, we identify that 1D RoPE is ill-suited for visual spatiotemporal correlation modeling, and while demonstrated to be useful, naive 3D RoPE exhibits imbalanced frequency spectra. Therefore, we propose MM‑RoPE, which preserves the original textual RoPE while seamlessly accommodating video data with comprehensive frequency spectra and scaled 3D positions. Secondly, to fit the video data’s nature and overcome the inefficiency of next-token decoding, we adopt a parallel and mask-based discrete diffusion with the intra-frame bidirectional and inter-frame causal attention masks. Based on this attention mask, we uncover the frame‑wise loss imbalance issue caused by spatial information redundancy and propose Autoregressive Discrete Diffusion Forcing, which introduces temporal tube masking during training with a compatible inference‑time masking policy to avoid quality degradation. Despite using only 48 GPUs for pre-training and fine-tuning, limited data and a discrete tokenizer, Lumos-1 achieves results surpassing those of Show-o2 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V.
Code and models: https://github.com/alibaba-damo-academy/Lumos
Email: hj.yuan@zju.edu.cn
Abstract
Autoregressive large language models (LLMs) have unified a vast range of language tasks, inspiring preliminary efforts in autoregressive (AR) video generation. Existing AR video generators either diverge from standard LLM architectures, depend on bulky external text encoders, or incur prohibitive latency due to next-token decoding. In this paper, we introduce Lumos-1, an LLM-based unified model for AR video generation with efficient discrete diffusion. Firstly, to fit videos with LLMs, we identify that 1D RoPE is ill-suited for visual spatiotemporal correlation modeling, and while demonstrated to be useful, naive 3D RoPE exhibits imbalanced frequency spectra. Therefore, we propose MM‑RoPE, which preserves the original textual RoPE while seamlessly accommodating video data with comprehensive frequency spectra and scaled 3D positions. Secondly, to fit the video data’s nature and overcome the inefficiency of next-token decoding, we adopt a parallel and mask-based discrete diffusion with the intra-frame bidirectional and inter-frame causal attention masks. Based on this attention mask, we uncover the frame‑wise loss imbalance issue caused by spatial information redundancy and propose Autoregressive Discrete Diffusion Forcing, which introduces temporal tube masking during training with a compatible inference‑time masking policy to avoid quality degradation. Despite using only 48 GPUs for pre-training and fine-tuning, limited data and a discrete tokenizer, Lumos-1 achieves results surpassing those of Show-o2 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V.
Code and models: https://github.com/alibaba-damo-academy/Lumos
Email: hj.yuan@zju.edu.cn
@@ 1 Introduction
Autoregressive (AR) models have demonstrated significant advancement in the field of language processing [48, 74, 49, 4, 42] by unifying diverse language tasks into one single framework and expanding the scale of large language models to unprecedented sizes. Following this trail of research, significant efforts have emerged in autoregressive visual generation [36, 7, 8, 46]. By utilizing similar architectural and generation designs as those used in LLMs, there is substantial potential for advancing LLMs towards a unified model capable of both visual generation and understanding [81, 86, 103, 72].
@@
Pioneering research efforts have been dedicated to instantiating AR visual generation, with a focus on key aspects, spanning improving autoregressive paradigm [50, 71, 8, 46], improving tokenizer format [36] (e.g., from discrete to continuous), and enhancing tokenizer capabilities [67, 78, 61]. However, an AR video generation paradigm that is fully compatible with LLM-based unified models (e.g., Chameleon [69] and EMU3 [81]) remains underexplored. Preliminary attempts either exhibit architectural distinctions with LLM architectures (e.g., NOVA [15] and Phenaki [76]), rely on external text encoders (e.g., LlamaGen [65] and Fluid [17]), or trail in generation efficiency due to next-token prediction (e.g., Loong [83]). Therefore, the introduction of LLMs to visual generation motivates careful designs for modeling the spatiotemporal visual space.
Adapting LLMs for video generation presents two fundamental challenges. First, the standard positional encoding in LLMs, such as 1D Rotary Position Embeddings (RoPE) [64], is inherently designed for sequential text and is ill-suited for modeling the complex 3D spatiotemporal correlations of video. Second, the standard autoregressive paradigm of next-token prediction, while effective for text, is notoriously inefficient for visual data and fails to properly model videos’ unique properties: spatial bidirectionality within a frame and strong temporal causality across frames. Addressing these two challenges—RoPE representation and prediction paradigm—is critical for building an effective and efficient model for video generation.
Table 1: Design comparison with other types of RoPE.
RoPE Type Compatiable with Text RoPE 3D Structure Comprehensive Frequency Allocation Strategic Scaling
M-RoPE
U-RoPE
IL-RoPE
VideoRoPE
HoPE
$M$-RoPE
@@
To account for the spatiotemporal nature of videos using the RoPE technique, we start with systematic experiments and analysis (in Sec.˜3.1) of extending 1D RoPE to a vanilla 3D RoPE [27, 33], revealing that 1D RoPE is far from optimal, and the incorporation of 3D RoPE facilitates autoregressive generative learning. However, existing 3D RoPE still suffers from imbalanced frequency spectrum ranges for temporal and spatial modeling, as shown in Sec.˜3.1. Building on this insight, we propose MM-RoPE, a new family of RoPE that better accommodates the structure of video data. As summarized in Tab.˜1, the core enhancement lies in more distributed frequency allocations for more comprehensive context modeling, substantially improving the visual generation capability. Moreover, we design a principled scaling to the 3D positions, which improves vision-language modality balancing. Though MM-RoPE modifies RoPE for visual tokens, the RoPE for text tokens remains consistent with native LLMs, therefore preserving the language learning capabilities.
@@
To account for the nature of videos (i.e., spatial bidirectionality and temporal causality), we build our model upon the intra-frame bidirectional and inter-frame causal token dependency strategy, so that we can introduce the use of the efficient and parallel discrete diffusion [8]. Discrete diffusion requires training with masks [36]. However, the naive global random mask prediction leads to the loss imbalance issue, which means that masked visual tokens in the later frames tend to have much lower loss, impeding effective video modeling. We attribute this to videos’ spatial information redundancy, as spatial information leakage allows the model to easily predict the masked token by attending to unmasked ones in previous frames. To address this, we propose Autoregressive Discrete Diffusion Forcing (AR-DF). The core of AR-DF training involves temporal tube masking [73], which repeats a frame-level random mask pattern across the temporal axis. Such a design prevents masked tokens in later frames from learning a shortcut–copying previous ones. The core of AR-DF inference is the strategic inference-time masks that align with partial observation of history during training, enabling inference without degraded frame quality and motion.
@@
Collectively, we propose Lumos-1, a model building on Llama [75] and discrete diffusion to achieve AR video generation. Through a stage-wise training with GPU memory-friendly techniques, we pre-train Lumos-1 from scratch on 60 million images and 10 million videos using only 48 GPUs, achieving performance comparable to EMU3 on GenEval, COSMOS-Video2World on VBench-I2V, and OpenSoraPlan on VBench-T2V (figure˜1 visualizes examples generated by Lumos-1).
Our contributions can be summarized as: 1) We propose Lumos-1, a pure LLM-based unified architecture with discrete diffusion for AR video generation. 2) We identify that although being useful, naive 3D RoPE exhibits imbalanced frequency spectra. Therefore, we propose MM-RoPE with more distributed frequency allocations and principled scaling, to achieve better dependency modeling and modality balancing. 3) We identify the loss imbalance issue during mask prediction training, and propose AR-DF to enforce training with temporal modeling and introduce inference-time masks to avoid quality degradation. 4) We validate the efficacy of the above designs through extensive experiments. Lumos-1 achieves competitive results on T2I, I2V and T2V tasks, while using limited training computing, data and a discrete tokenizer.
@@ 2 Related Work
Autoregressive video generation. Autoregressive video synthesis follows a trajectory similar to images (from a dedicated architecture to a unified one like LLMs), but the substantial computational overhead encouraged researchers to explore different granularities of autoregression. At the macro level, methods such as CausVid [92], Pyramidal Flow [31], and MAGI-1 [60] generate video clips by recursively invoking diffusion models. A second line of work adopts a hybrid AR-diffusion strategy that couples a pre-trained diffusion backbone with an external autoregressive planner, as exemplified by ARLON [38] and Mardini [44]. Our focus, however, is on micro AR approaches [22, 104, 96, 97, 76, 32, 81, 15, 83], which treat the entire spatiotemporal token sequence as a single context and generate videos end-to-end with one autoregressive transformer (e.g., Phenaki [76], VideoPoet [32], Loong [83], EMU3 [81], NOVA [15]). Building on this paradigm, we systematically study the use of RoPE, and further introduce AR-DF, which ensures efficiency and is compatible with intra-frame bidirectionality and inter-frame temporal causality, enabling effective training and inference with LLM architectures.
RoPE for vision-language data. RoPE was first proven effective for LLMs [64, 74, 28] and later adopted in DiTs [51] to inject spatial priors. However, its potential for AR video generation is heavily underexplored. Related works, including M-RoPE [80], U-RoPE [68], RoPE-2D [2], VoPE [10], TAD-RoPE [18], IL-RoPE [39], VideoRoPE [84] and HoPE [35], incorporated structure priors into RoPE (i.e., spatiotemporal information or its subsets) or improve the frequency allocation of RoPE for better vision-language understanding. However, they bear drawbacks: 1) some of them are not compatible with the original text RoPE; 2) their frequency allocations remain suboptimal for local-global dependency modeling; 3) their scaling strategy remains suboptimal for spatiotemporal data. We propose MM‑RoPE, a distributed and scaled RoPE technique that improves AR video synthesis while remaining plug‑and‑play for unified models.
@@ 3.1 Spatiotemporal Correlation Injection via MM-RoPE
Preliminaries of 3D RoPE. RoPE [64] aims to encode the absolute position with a rotation matrix while incorporating the explicit relative position dependency in the attention mechanism. If we denote $f_{q}(\bm{x}_{m},m)$ and $f_{k}(\bm{x}_{n},n)$ as query and key features encoding positions $m$ and $n$, the attention calculation in RoPE can be rewritten as:
$f_{q}(\bm{x}_{m},m)^{\rm T}f_{k}(\bm{x}_{n},n)=\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}\bm{R}^{d}_{\Theta,\tau}\bm{W}_{k}\bm{x}_{n},\quad\tau=n-m$ (1)
where $\bm{W}_{q,k}$ is the projection matrix; $\bm{R}^{d}_{\Theta,\tau}$ is the rotary matrix with pre-defined parameters $\Theta=\{\theta_{i}=\beta^{-2(i-1)/d},i=[1,2,...,d/2]\}$ ($d$ is the feature dimension and $\beta$ is the base frequency). We formulate $\bm{R}^{d}_{\Theta,\tau}$ using a base rotary matrix $R_{\theta,\tau}$, with $\theta$ as the frequency and $\tau$ as the relative position:
$\scriptsize\bm{R}^{d}_{\Theta,\tau}=\begin{bmatrix}R_{\theta_{1},\tau}&0&\cdots&0\\ 0&R_{\theta_{2},\tau}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\cdots&R_{\theta_{d/2},\tau}\\ \end{bmatrix},\quad R_{\theta,\tau}=\begin{bmatrix}\cos\tau\theta&-\sin\tau\theta\\ \sin\tau\theta&\cos\tau\theta\end{bmatrix}$} (2)
However, applying the original RoPE to modeling visual data remains suboptimal considering the spatiotemporal nature of visual tokens. Diffusion models [26] improved upon this by proposing 3D RoPE that jointly injects spatiotemporal latent coordinates during attention calculation [27, 33]. If we slightly abuse the annotation by denoting $\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}$ and $\bm{W}_{k}\bm{x}_{n}$ as $\bm{X}_{m}^{\rm T}$ and $\bm{X}_{n}$, we can write the attention calculation based on 3D RoPE as:
$\begin{aligned} \bm{X}_{m,t_{s}:t_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{t_{s}+1},\tau_{t}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{t_{e}},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,t_{s}:t_{e}}+\bm{X}_{m,h_{s}:h_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{h_{s}+1},\tau_{h}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{h_{e}},\tau_{h}}\\ \end{bmatrix}\bm{X}_{n,h_{s}:h_{e}}+\bm{X}_{m,w_{s}:w_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{w_{s}+1},\tau_{w}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{w_{e}},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,w_{s}:w_{e}}\end{aligned}$
where $\{t_{s},t_{e}\}=\{0,\frac{2}{16}d\}$, $\{h_{s},h_{e}\}=\{\frac{2}{16}d,\frac{5}{16}d\}$ and $\{w_{s},w_{e}\}=\{\frac{5}{16}d,\frac{1}{2}d\}$ denote the start and end dimension index for encoding temporal, height and width relative position; $\bm{X}_{m,t_{s}:t_{e}}^{\rm T}$ denotes the submatrix extracted from $\bm{X}_{m}^{\rm T}$ using row indices $2t_{s},2t_{e})$; other matrices are similarly defined.
@@
Initial exploration of incorporating 3D RoPE. Our preliminary exploration starts with the introduction of 3D RoPE to AR video generation. We follow previous works [17, 36] to utilize the validation loss to observe the effectiveness due to its strong correlation with evaluation metrics. By default, we use the cross-entropy loss (C-Loss), following standard LLM training. We compare the vanilla LLM RoPE with three schemes, as shown in figure˜2: 1) Scheme 1, which allocates the first $1/2$ channels to encode the global position (i.e., the index in the global sequence) and the second $1/2$ channels to encode the temporal, height and width positions with a ratio of $2:3:3$. For text tokens, we only use the first half channels to encode the global positions to ensure the language modeling capability, while for visual tokens, we only use the second half to encode 3D positions; 2) Scheme 2, which extends upon scheme 1 by leveraging the first half channels of the visual tokens to encode the global positions. 3) M-RoPE [80], which uses all channels of the visual tokens to encode the 3D positions. From figure˜3(a), we can observe that: 1) The incorporation of the spatiotemporal correlation in RoPE can significantly improve the performance of the model fitting data by comparing vanilla RoPE and Scheme 1. 2) The injection of the raster-scan order position information to visual tokens (i.e., the global position in Scheme 2 of figure˜2(a)) can degrade the performance. 3) A comprehensive channel utilization (M-RoPE) is better than partial channel utilization (Scheme 1). Therefore, it is promising to inject such priors in Lumos-1.
@@
Peering into 3D RoPE and identifying its limitations. Although 3D RoPE proves effective in practice, its design remains suboptimal. In figure˜3(b), we visualize how the frequencies are allocated to model temporal, height, and width dimensions. We observe that the temporal channels dominate large frequency spectrum, whereas the height and width channels are relegated to near-zero frequencies. For the sine function, the relative positions $\tau$ (when $\tau>=0$) should not exceed one period to avoid ambiguity since radians beyond $2\pi$ start repeating patterns in the function. Beyond this range, the model cannot distinguish fine-grained positional differences since rotary embeddings of this channel lose uniqueness. For the low-indexed channels, the embeddings can rotate significantly faster than high-indexed channels (as shown in figure˜3(c)), leading to accelerated aliasing and loss of embedding uniqueness. For the high-indexed channels, the embeddings rotate so slowly that they lack sufficient resolution to model subtle local changes. Finally, while height and width are symmetrically important, they occupy disproportionately small and different segments of the overall frequency spectrum, reducing their capacity to capture spatial details effectively.
@@
MM-RoPE: a distributed and scaled 3D RoPE mechanism. To elegantly solve the aforementioned limitation, we propose MM-RoPE, a distributed 3D RoPE mechanism. Compared with M-RoPE [80] that is widely adopted in vision-language models, one core idea of MM-RoPE is to encode relative positions across comprehensive frequency spectra for all 3D information. As illustrated in figure˜2(b), the RoPE for text tokens in MM-RoPE follows the standard design of an LLM, whereas the RoPE for visual tokens is comprised of multiple meta MM-RoPE components. Within each meta MM-RoPE, we keep the ratio of 3D information identical to 3D RoPE (i.e., $2:3:3$), while minimizing the overall dimensionality to maintain a more distributed design. Concretely, we first allocate channels for temporal modeling, then symmetrically interleave height and width channels to model spatial information. We can formulate the attention calculation of the first meta MM-RoPE as:
$\bm{X}_{m,0:2}^{\rm T}\begin{bmatrix}R_{\theta_{1},\tau_{t}}&0\\ 0&R_{\theta_{2},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,0:2}+\bm{X}_{m,2:4}^{\rm T}\begin{bmatrix}R_{\theta_{3},\tau_{h}}&0\\ 0&R_{\theta_{4},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,2:4}$ $+$ (4)
$\bm{X}_{m,4:6}^{\rm T}\begin{bmatrix}R_{\theta_{5},\tau_{h}}&0\\ 0&R_{\theta_{6},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,4:6}+\bm{X}_{m,6:8}^{\rm T}\begin{bmatrix}R_{\theta_{7},\tau_{h}}&0\\ 0&R_{\theta_{8},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,6:8}$
@@
3:  $\mathcal{C}\leftarrow G_{\phi}(\bm{X}_{p},AttnMask^{(p)})$ $\triangleright$ Store cache for the prompt